OpenStudy (anonymous):

Sin-¹(-1/2) without the inverse function

OpenStudy (anonymous):

\[Sin^{-1} (\frac{ -1 }{ 2 })\]

OpenStudy (anonymous):

\(\large y=Sin^{-1}(\frac{-1}{2}) \rightarrow siny=\frac{-1}{2} \) now can you determine y?

OpenStudy (anonymous):

oh just move sides

OpenStudy (anonymous):

remember the domain of arcsin....

OpenStudy (anonymous):

oh yes !

OpenStudy (anonymous):

y = ???

OpenStudy (anonymous):

just remember that because that is a capital Sin^(-1), that means "principle"

OpenStudy (anonymous):

how would i find the exact value of that @dpaInc

OpenStudy (anonymous):

sorry i haven't learned this sorry

OpenStudy (anonymous):

anyone know ? :$

OpenStudy (anonymous):

immm really sorry